The body, weighing 4 kg, is released from a height of 30 cm above the spring. The coefficient of elasticity of the spring is 1600 N / m. Find how much will be the biggest compression of the spring?


1
Expert's answer
2021-05-18T11:11:10-0400

As the body lands on the spring, the compression can be found by the conservation of energy: the gravitational potential energy is wasted to compress the spring. The gravitational potential energy is


"GPE=mg(h+x),"

the elastic potential energy is


"EPE=\frac12 kx^2,\\\space\\
GPE=EPE.\\\space\\
mg(h+x)=\frac12 kx^2,\\\space\\
\frac12kx^2-mgx-mgh=0,\\\space\\
x=\frac{mg+\sqrt{(mg)^2+2kmgh}}{k}=0.15\text{ m}."